OpenStudy (anonymous):

how to solve (log 24)(log x^2)=5.log 12

OpenStudy (unklerhaukus):

is this your equation?\[\log(24)\cdot\log(x^2)=5\cdot\log(12)\]

OpenStudy (unklerhaukus):

and you have to solve for \(x\)?

OpenStudy (unklerhaukus):

... @seshani, are you still here?

OpenStudy (unklerhaukus):

?

OpenStudy (unklerhaukus):

(i need to be sure i have interpreted the question correctly, Before i try to guide to to solve it)

OpenStudy (anonymous):

(log 24) (log x^2) = 5.log 12 . this is the question